OpenStudy (anonymous):

The y-intercept of the line whose equation is 2x - 3y = 6 is -2??

OpenStudy (anonymous):

looks good. Why do u need to check:-)

OpenStudy (anonymous):

y =-2

OpenStudy (anonymous):

hey you changed the question

OpenStudy (anonymous):

doesn't trust us !!!?

OpenStudy (amistre64):

one rather trivial way to determine intercepts in a line equation is to divide it all by that constant. 2x - 3y = 6 ; /6 2x -3y 6 --- + --- = --- ; and simplify 6 6 6 x -y -- + --- =1 ; and if theres a negative place it on the bottom 3 2 for easy pickings \[\frac{x}{3}+\frac{y}{-2}=1\] now the intercepts can be read by the bottom numbers

OpenStudy (anonymous):

y=mx+b form yields \[2x-6=3y\] --> \[y=2/3*x-2\] y-intercept is the value of the equation when x=0, therefore you are right